# Latent space models for grouped multiplex networks

### v1

@@ 3.1 Likelihood maximization with nuclear norm penalization
In Section 4, we will derive theoretical guarantees for the estimators produced by Algorithm 1 under the assumption of Gaussian edge entry distribution and unit learning rate $\eta_{1}=\eta_{2}=1$. In this case, the proximal gradient descent updates are especially convenient for theoretical analysis. For the $k$-th problem in the first stage, the update rule at iteration $t\geq 1$ obtains the form
$R_{k\ell}^{(t)}=\mathcal{T}_{\lambda_{1k}\alpha_{1k\ell}}\Bigl[A_{k\ell}-{(S+Q_{k})}^{(t-1)}\Bigr],\qquad\ell=1,\ldots,m_{k},$ (16)
$(S+Q_{k})^{(t)}=\mathcal{T}_{\lambda_{1k}/m_{k}}\Bigl[{1\over m_{k}}\sum_{\ell=1}^{m_{k}}\Bigl(A_{k\ell}-R_{k\ell}^{(t)}\Bigr)\Bigr].$
Similarly, the second stage updates can be written for $t\geq 1$ as
$Q_{k}^{(t)}=\mathcal{T}_{\lambda_{2}\alpha_{2k}/m_{k}}\Bigl[{1\over m_{k}}\sum_{\ell=1}^{m_{k}}\Bigl(A_{k\ell}-\hat{R}_{k\ell}-S^{(t-1)}\Bigr)\Bigr],\qquad k=1,\ldots,K,$ (17)
$S^{(t)}=\mathcal{T}_{\lambda_{2}/M}\Bigl[{1\over M}\sum_{(k,\ell)\in\mathcal{I}}\Bigl(A_{k\ell}-\hat{R}_{k\ell}-Q_{k}^{(t)}\Bigr)\Bigr].$
Notice that with unit learning rates, the update rule for each parameter does not depend on its value in the previous iteration. While these properties help us simplify our theoretical analysis, tuning of the learning rate is highly recommended in practice to achieve a reasonable convergence speed.
3.2 Hyperparameter tuning
In this section, we propose a procedure for tuning the hyperparameters $\{\lambda_{1k}\}_{k=1}^{K},\lambda_{2},$ $\{\alpha_{2k}\}_{k=1}^{K},$ and $\{\alpha_{1k\ell}\}_{\mathcal{I}}$ used in Algorithm 1. Our approach is inspired by the commonly used edge cross-validation method of li_network_2020.
At first glance, the total number of hyperparameters that need to be tuned in Algorithm 1 may seem overwhelming. However, there are several important considerations that significantly simplify our cross validation procedure. Note that $\lambda_{1k}$ with $\{\alpha_{1k\ell}\}_{\ell=1}^{m_{k}}$ only appear in (10) and $\lambda_{2}$ with $\{\alpha_{2k\ell}\}_{\ell=1}^{m_{k}}$ only in (11), which means we can tune these groups of hyperparameter separately. Since each group is still too large for tuning, we can fix $\alpha_{1k}:=\alpha_{1k\ell}$ for $(k,\ell)\in\mathcal{I}$ and $\alpha_{2}:=\alpha_{2k}$ for $k=1,\ldots,K$. Then, it is enough to tune only two parameters in each of the subproblems. As a simpler alternative, it is also possible to optimize only the more influential $\lambda_{2},\{\lambda_{1k}\}_{k=1}^{K}$ parameters appearing in all the nuclear norm penalties of the corresponding problems. In this case, $\{\alpha_{1k\ell}\}_{\mathcal{I}}$ and $\{\alpha_{2k}\}_{k=1}^{K}$ are assigned following the theoretically optimal values derived for the Gaussian edge entry distribution in Corollary 4.1. In our simulations, this approach produced errors comparable to the more expensive tuning of two hyperparameters per subproblem, so we use it as the default option in our implementation.
To tune the hyperparameters of the $k$-th subproblem in the first stage, we propose sample a random subset of “training” edge entries in layers of the $k$-th group:
$\mathcal{A}^{(k)}_{train}\subset\mathcal{A}^{(k)}:=\{A_{k\ell,ij}:\ell=1,\ldots,m_{k},\ 1\leq i\leq j\leq n\},$
solve Problem (10) with log-likelihood terms in $\mathcal{L}_{k}$ restricted to edge entries in $\mathcal{A}^{(k)}_{train}$ and then evaluate the non-penalized likelihood (9) on the remaining “test” entries $\mathcal{A}_{test}^{(k)}=\mathcal{A}^{(k)}\setminus\mathcal{A}^{(k)}_{train}$. This procedure is repeated for each combination of hypeparameters in the predefined grid, with the test log-likelihood possibly averaged across multiple cross-validation folds for stability. Similarly, for tuning $\lambda_{2}$ and $\alpha_{2}:=\alpha_{2k},\ k=1,\ldots,K$ in Problem (11), we can randomly sample a subset of edge entries from all the layers
$\mathcal{A}_{train}\subset\mathcal{A}:=\{A_{k\ell,ij}:k=1,\ldots,K,\ \ell=1,\ldots,m_{k},\ 1\leq i\leq j\leq n\},$
solve the problem with log-likelihood terms restricted to $\mathcal{A}_{train}$, and evaluate the non-penalized log-likelihood (8) on $\mathcal{A}_{test}=\mathcal{A}\setminus\mathcal{A}_{train}$.
3.3 Parameter initialization
In this section, we discuss different ways to define the initialization procedures $\it{FirstStageInit}$ and $\it{SecondStageInit}$ in Algorithm 1. Due to the convexity of the problems (10) and (11), we expect initialization to impact only the number of iterations needed for convergence but not the algorithm’s attaining the global optimum. From the perspective of our theoretical framework, initialization is more crucial as our consistency result relies on the initializer being sufficiently close to the ground truth.
For the Gaussian edge-entry model, we propose to initialize the shared component of a collection of layers as their average and the individual components as the difference between the layer and the shared component initializer. When the ranks of the latent components are known or preliminarily estimated, we can also truncate the initializers, as defined in (15). In our case, due to the first identifiability assumption in Proposition 2.1, the rank of the total shared component for layers in group $k$ is $d_{0}+d_{k}$, implying the initializer:
For exponential family distributions with a non-linear link function $g$, we propose first getting a proxy of $\Theta_{k\ell}$ by apply $g^{-1}$ to each layer $A_{k\ell}$ element-wise, then truncate the result to [5,5][-5,5], and then average the corresponding transformations instead of the layers themselves.
Alternatively, one can use the recently developed “shared space hunting” (SSH) approach (see Algorithm 1 in TianYinqiu2024). In contrast to the averaging initialization, this method is computationally more expensive as it requires first estimating the latent positions $X_{k\ell}$ of the nodes in each layer and thus their latent ranks $D_{k\ell}$ in particular. Moreover, our simulations (not presented in this work, but available on GitHub) suggest that the SSH can produce very poor initializers if the latent dimensions are inaccurately estimated, which turned out to be the case for small $n$ and large ground-truth ranks. On the other hand, even when $n$ is large and ranks are small, the averaging approach appears to produce only marginally worse results than SSH. So due to its robustness and low computational cost, we use the averaging initializer as our default option.
For the $\it{SecondStageInit}$, we have more options to choose from since it can use the first-stage estimates $\{\widehat{S+Q_{k}}\}_{k=1}^{K}$ and $\{\hat{R}_{k\ell}\}_{\mathcal{I}}$ besides the layers $\{A_{k\ell}\}_{\mathcal{I}}$ themselves. The natural candidates for the initializer of the shared component $S$ would be to average the estimates $\widehat{S+Q_{k}}$
or the group-wise mean residuals between the layer and the individual component estimate
Similarly to the first-stage initializer in (18), each layer $A_{k\ell}$ in (20) has to be preliminary transformed by the inverse link function $g^{-1}$ and then truncated for non-Gaussian distributions. Alternatively, the SSH approach can be used for the initial estimation of the shared component $S$ from either $\{\widehat{S+Q}_{k}\}_{k=1}^{K}$ or $\Bigl\{{1\over m_{k}}\sum_{\ell=1}^{m_{k}}(A_{k\ell}-\hat{R}_{k\ell})\Bigr\}_{k=1}^{K}$.
In our implementation of Algorithm 1, we used (20) as the second-stage initializer, as it draws a more direct parallel to the MultiNeSS fitting procedure per Proposition 3.1. However, empirically, we observed that both (20) and (19) imply almost indistinguishable convergence speeds for Problem (11). To give a possible theoretical explanation to this observation, we derive an explicit relationship between the corresponding terms in the two initializing averages under the Gaussian edge entry distribution and a mild extra assumption on the self-loop distribution. In particular, the following proposition demonstrates that in the Gaussian case, $\widehat{S+Q_{k}}$ is the soft-thresholded version of the averaged residuals ${1\over m_{k}}\sum_{\ell=1}^{m_{k}}\bigl(A_{k\ell}-\hat{R}_{k\ell}\bigr)$. The proof of this proposition can be found in Section B.2 of the Appendix.

### v2

@@ 3.1 Likelihood maximization with a nuclear norm penalty
3.2 Choice of tuning parameters
In this section, we propose a procedure for choosing the hyperparameters $\{\lambda_{1k}\}_{k=1}^{K},\lambda_{2},$ $\{\alpha_{2k}\}_{k=1}^{K},$ and $\{\alpha_{1k\ell}\}_{\mathcal{I}}$ used in Algorithm 1, based on the commonly used edge cross-validation method of Li et al. (2020).
At first glance, the total number of tuning parameters seems overwhelming, but there are several ways to significantly simplify cross-validation. Note that $\lambda_{1k}$ with $\{\alpha_{1k\ell}\}_{\ell=1}^{m_{k}}$ only appear in (10) and $\lambda_{2}$ with $\{\alpha_{2k\ell}\}_{\ell=1}^{m_{k}}$ only in (11), which means we can tune these groups of parameters separately. We further fix $\alpha_{1k}:=\alpha_{1k\ell}$ for $(k,\ell)\in\mathcal{I}$ and $\alpha_{2}:=\alpha_{2k}$ for $k=1,\ldots,K$, resulting in only two parameters to tune for each of the subproblems. As a simpler alternative, one could also just tune the more consequential $\lambda_{2},\{\lambda_{1k}\}_{k=1}^{K}$ parameters, and set $\{\alpha_{1k\ell}\}_{\mathcal{I}}$ and $\{\alpha_{2k}\}_{k=1}^{K}$ to their theoretically optimal values derived for the Gaussian edge entry distribution in Corollary 4.1. In our simulations, this approach gave results comparable to tuning two hyperparameters per subproblem, and we set it as the default option in our implementation.
To tune the hyperparameters of the $k$-th subproblem in the first stage, we sample a random subset of “training” node pairs in layers of the $k$-th group:
$\mathcal{A}^{(k)}_{train}\subset\mathcal{A}^{(k)}:=\{A_{k\ell,ij}:\ell=1,\ldots,m_{k},\ 1\leq i\leq j\leq n\}.$
We then solve (10) with log-likelihood terms in $\mathcal{L}_{k}$ restricted to node pairs in $\mathcal{A}^{(k)}_{train}$ and then evaluate the non-penalized likelihood (9) on the remaining “test” node pairs $\mathcal{A}_{test}^{(k)}=\mathcal{A}^{(k)}\setminus\mathcal{A}^{(k)}_{train}$, and choose the parameters from a pre-defined grid to optimize the test log-likelihood averaged across multiple cross-validation folds for stability. Hyperparameters in the second stage problem (11) are tuned similarly, with the only difference that node pairs are sampled from $A_{k\ell},(k,\ell)\in\mathcal{I}$.
3.3 Initialization
Both stages of the Algorithm 1 require initial values, but they are convex problems, so the choice of initialization primarily impacts how long it takes the algorithm to converge, not its ability to attain the global optimum. However, our theoretical results rely on the initializer being sufficiently close to the truth.
For the Gaussian edge-entry model, we initialize the shared component of a collection of layers as their average and the individual components as the resulting residuals. When the ranks are known, these can be further truncated, resulting in initializers
For exponential family distributions with a non-linear link function $g$, we first get a proxy of $\Theta_{k\ell}$ by applying $g^{-1}$ to each layer $A_{k\ell}$ element-wise (truncating if necessary to avoid the inverse going to infinity), and then average the transformed layers.
Alternatively, one can use the recently developed “shared space hunting” (SSH) approach (Algorithm 1 in Tian et al. (2024)). This method is computationally more expensive than averaging as it requires first estimating the latent positions $X_{k\ell}$ of the nodes in each layer and the latent space dimensions $D_{k\ell}$. Our simulations (available on GitHub) suggest that the SSH can produce very poor initializers if the latent dimensions are inaccurately estimated, which tends to happen when $n$ is small and/or the true ranks are large, whereas with correctly estimated ranks the SSH approach produces only marginally better results than averaging. Thus we use averaging as the default initialization option, due to its robustness and low computational cost.
For Stage II, we have more initialization options to choose from since it can use the first-stage estimates $\{\widehat{S+Q_{k}}\}_{k=1}^{K}$ and $\{\hat{R}_{k\ell}\}_{\mathcal{I}}$. The natural candidates for the initializer of the shared component $S$ would be the average of either the estimates $\widehat{S+Q_{k}}$,
or of the residuals between the adjacency matrix and the individual component estimate,
Similarly to the first-stage initializer in (15), for non-Gaussian distributions each layer $A_{k\ell}$ in (17) should be first transformed by the inverse link function $g^{-1}$ and then truncated. Alternatively, the SSH approach can be used for the initial estimation of the shared component $S$ from either $\{\widehat{S+Q}_{k}\}_{k=1}^{K}$ or $\Bigl\{{1\over m_{k}}\sum_{\ell=1}^{m_{k}}(A_{k\ell}-\hat{R}_{k\ell})\Bigr\}_{k=1}^{K}$.
In our implementation of Algorithm 1, we used (17) as the second-stage initializer, as it gives the most direct parallel to MultiNeSS per Proposition 3.1. However, empirically we observed that both initializations (17) and (16) result in essentially the same convergence speeds for problem (11). To provide a possible theoretical explanation to this observation, we derive an explicit relationship between the two under the Gaussian edge distribution and a mild extra assumption on the self-loop distribution. The following proposition demonstrates that in the Gaussian case, $\widehat{S+Q_{k}}$ is the soft-thresholded version of the averaged residuals ${1\over m_{k}}\sum_{\ell=1}^{m_{k}}\bigl(A_{k\ell}-\hat{R}_{k\ell}\bigr)$. The proof can be found in Section B.2 of the Appendix.
Proposition 3.2.
Assume that the edge entry distribution is $f(\cdot;\theta,\sigma)=\mathcal{N}(\theta,\sigma^{2})$ for the layers’ off-diagonal entries and $\mathcal{N}(\theta,2\sigma^{2})$ for the diagonal entries. Then, the estimates produced by solving problem (10) are related as follows:
4 Consistency results
This section establishes theoretical guarantees for the parameter estimates obtained by running Algorithm 1 with unit learning rates $\eta_{1}=\eta_{2}=1$ under the Gaussian assumption on edge distribution, $f(\cdot;\theta,\sigma)=\mathcal{N}(\theta,\sigma^{2})$. To simplify our analysis, we assume that Stage II and each subproblem of Stage I are terminated after the first iteration $(t=1)$. Throughout this section, we assume that $M$, $K$, $m_{k}$, $d_{0}$, $d_{k}$, and $d_{k\ell}$ are possibly increasing functions of $n$. The variance $\sigma^{2}$ is assumed fixed. All proofs for this section can be found in Section B.3 of the Appendix.
We begin by introducing additional notation. For $a,b\in\mathbb{R}$, denote $a\vee b=\max(a,b)$. For real-valued sequences $g_{n},h_{n}$, we write $g_{n}\lesssim h_{n}$ if $g_{n}=O(h_{n})$ and $g_{n}\asymp h_{n}$ if $g_{n}\lesssim h_{n}$ and $h_{n}\lesssim g_{n}$. For brevity, we write
Note there is a one-to-one correspondence between these newly defined thresholds and the original parameters $\{\lambda_{1k}\}_{k=1}^{K},\lambda_{2},\{\alpha_{2k}\}_{k=1}^{K},\{\alpha_{1k\ell}\}_{\mathcal{I}}$. For matrices $Z_{1},Z_{2}\in\mathbb{R}^{n\times n}$ we denote the cosine similarity between their eigenspaces by
For convenience, we also define maximal angles between the following pairs of latent component types
$s_{v,w}=\max_{1\leq k\leq K}\cos(S,Q_{k}),$ $s_{w,w}$ $=\max_{1\leq k_{1}<k_{2}\leq K}\cos(Q_{k_{1}},Q_{k_{2}}),$
$s_{vw,u}^{(k)}=\max_{1\leq\ell\leq m_{k}}\cos(S+Q_{k},R_{k\ell}),$ $s_{u,u}^{(k)}$ $=\max_{1\leq\ell_{1}<\ell_{2}\leq m_{k}}\cos(R_{k\ell_{1}},R_{k\ell_{2}}),$
$s_{u,u}=\max_{(k_{1},\ell_{1})\neq(k_{2},\ell_{2})}\cos(R_{k_{1}\ell_{1}},R_{k_{2}\ell_{2}}).$
These quantities will appear in the error bounds on the latent components, supporting the intuition that the accuracy of separating latent spaces depends on how similar they are to each other.
Before we proceed to the main result, we state a concentration bound that depends on the Gaussian assumption. With a minor modification of the algorithm, the main result can be extended to sub-Gaussian distributions, as discussed below in Remark 7. Rewriting the observed matrix as signal plus noise, we have
$A_{k\ell}=\Theta_{k\ell}+E_{k\ell}=(S+Q_{k}+R_{k\ell})+E_{k\ell},$ (19)
where $E_{k\ell}\in\mathbb{R}^{n\times n}$ is a symmetric centered noise matrix with $E_{k\ell,ij}\stackrel{{\scriptstyle\text{ind}}}{{\sim}}\mathcal{N}(0,\sigma^{2})$ for $1\leq i\leq j\leq n$ and $(k,\ell)\in\mathcal{I}$, Gaussian matrices enjoy a convenient concentration bound on their operator norm, as well as the averages of their independent copies. In particular, our theoretical analysis will be restricted to the event where all individual errors $E_{k\ell},(k,\ell)\in\mathcal{I}$, their groupwise averages $\bar{E}_{k}={1\over m_{k}}\sum_{\ell=1}^{m_{k}}E_{k\ell},k=1,\ldots,K$, and total average $\bar{E}={1\over M}\sum_{(k,\ell)\in\mathcal{I}}E_{k\ell}$ are bounded as follows:
The following lemma shows that this event has a high probability. The proof is analogous to Lemma 3 in (MacDonald et al., 2022).
Lemma 4.1.
With a universal constant $C_{0}>0$, it holds $\mathbb{P}\bigl(\mathcal{E}_{noise}\bigr)\geq 1-(M+K+1)ne^{-C_{0}n}$.
Next, we state assumptions needed to establish consistency. The first says that all groups have comparable sizes, asymptotically dominating the number of groups.
Assumption 1.
For each group $k=1,\ldots,K$, $m_{k}\asymp M/K$ with $K\lesssim M^{1/2}$ and $M\rightarrow\infty$.
The next assumption controls the signal-to-noise ratio, calibrated relative to the concentration bound in Lemma 4.1.
Assumption 2.
There are constants $0<b_{S}<B_{S}$, $0<b_{Q}<B_{Q}$, $0<b_{S+Q}<B_{S+Q}$, $0<b_{R}<B_{R}$, and $\tau\in(1/2,1]$, such that
$b_{S}n^{\tau}\leq\left|\gamma_{d_{0}}\left(S\right)\right|\leq\left|\gamma_{1}\left(S\right)\right|\leq B_{S}n^{\tau},$ (21)
$b_{Q}n^{\tau}\leq\left|\gamma_{d_{k}}\left(Q_{k}\right)\right|\leq\left|\gamma_{1}\left(Q_{k}\right)\right|\leq B_{Q}n^{\tau},$ $k=1,\ldots,K,$
$b_{S+Q}n^{\tau}\leq\left|\gamma_{d_{0}+d_{k}}\left(S+Q_{k}\right)\right|\leq\left|\gamma_{1}\left(S+Q_{k}\right)\right|\leq B_{S+Q}n^{\tau},$ $k=1,\ldots,K,$
$b_{R}n^{\tau}\leq\left|\gamma_{d_{k\ell}}(R_{k\ell})\right|\leq\left|\gamma_{1}(R_{k\ell})\right|\leq B_{R}n^{\tau},$ $(k,\ell)\in\mathcal{I}$
Since we assumed that both Stage I and Stage II of Algorithm 1 are run for one iteration only, it is important to require that the initializers for each stage are "good enough". Rather than state the result for any particular choice of initializers, we formulate this requirement as a separate generic assumption. We will later show that our choice of initializer satisfies this assumption.
Assumption 3.
For $n$ sufficiently large, Stage I and II initializers on the set $\mathcal{E}_{noise}$ satisfy
$\mathbb{P}\Bigl\{\|S+Q_{k}-(S+Q_{k})^{(0)}\|_{2}\leq r_{k}^{(I)},\ \|S-S^{(0)}\|_{2}\leq r^{(II)}\,|\,\mathcal{E}_{noise}\}=1,$ (22)
where $r^{(I)}_{k}$ and $r^{(II)}$ are some $o(n^{\tau})$ deterministic functions of $n$.
In Proposition 4.1, we state explicit conditions for the initializers (15) and (17) to satisfy this assumption with errors of order $\sqrt{n}$. Alternatively, Theorem 1 of Tian et al. (2024) provides different conditions under which the Frobenius norm (and thus the operator norm) of the SSH initializer’s error is of order $\sqrt{n}\log n$.
With Assumptions 1, 2, and 3 at hand, we are ready to state the main result.
Theorem 4.1.
Suppose the edge-entry distribution is $f(\cdot;\theta,\sigma)=\mathcal{N}\left(\theta,\sigma^{2}\right)$ with fixed $\sigma^{2}$. Then under Assumptions 1, 2, 3, with probability greater than $1-(M+K+1)ne^{-C_{0}n}$, where $C_{0}>0$ is a universal constant, and for $n$ sufficiently large, the estimates produced by Algorithm 1 with learning rates $\eta_{1}=\eta_{2}=1$ and without refitting satisfy, for $(k,\ell)\in\mathcal{I}$,
$\begin{array}[]{lll}\textnormal{(Stage I)}&\|\hat{R}_{k\ell}-R_{k\ell}\|_{F}\leq 4d_{k\ell}^{1/2}\rho_{1k\ell},&\|\widehat{S+Q}_{k}-(S+Q_{k})\|_{F}\leq 4(d_{0}+d_{k})^{1/2}\rho_{1k}\\ \textnormal{(Stage II)}&\|\hat{Q}_{k}-Q_{k}\|_{F}\leq 4d_{k}^{1/2}\rho_{2k},&\|\hat{S}-S\|_{F}\leq 4d_{0}^{1/2}\rho_{2},\end{array}$
where
$\rho_{1k\ell}\asymp r_{k}^{(I)}\vee n^{1/2},$
$\rho_{1k}\asymp\max_{1\leq\ell\leq m_{k}}\rho_{1k\ell}\Bigl[{1\over m_{k}}\vee s_{u,u}^{(k)}\vee n^{-\tau}\max_{1\leq\ell\leq m_{k}}\rho_{1k\ell}\Bigr]^{1/2},$
$\rho_{2k}\asymp r^{(II)}\vee\rho_{1k},$
$\rho_{2}\asymp\max_{1\leq k\leq K}\rho_{2k}\Bigl[{1\over K}\vee s_{w,w}\vee n^{-\tau}\max_{1\leq k\leq K}\rho_{2k}\Bigr]^{1/2}\vee\max_{(k,\ell)\in\mathcal{I}}\rho_{1k\ell}$ $\Bigl[{1\over M}\vee s_{u,u}\vee n^{-\tau}\max_{(k,\ell)\in\mathcal{I}}\rho_{1k\ell}\Bigr]^{1/2},$
and the constants in the rates depend only on $(B_{R},b_{R},B_{S+Q},b_{S+Q},B_{Q},b_{Q},B_{S},b_{S},\sigma)$. Additionally, if $S$, $Q_{k},$ and $R_{k\ell}$ are positive semi-definite (PSD), then $\hat{S},\hat{Q}_{k},\hat{R}_{k\ell}$ are also PSD.
@@ 4 Consistency results
An important advantage of Algorithm 1 is the convexity of the optimization problems, which should mean the initial values do not affect consistency, only the speed of convergence. We made an assumption on the initializer error in Theorem 4.1 because we only analyze one update of the gradient descent, due to the complicated alternating updates. We leave it for future work to formally remove the initialization error assumption, though by convexity it is clear that even if the chosen initial value does not satisfy it (and we show that ours do), after some number of gradient descent steps it will.
Remark 7.
Our proof of Theorem 4.1 relies on two key properties of the Gaussian distribution: (i) the convenient form of the proximal gradient updates for the Gaussian log-likelihood, and (ii) the concentration result of Bandeira and van Handel (2016) used to establish Lemma 4.1. The concentration result was extended to sub-Gaussian distributions in the same paper (Corollary 3.3). The convenient form of gradient updates can be retained by replacing the log-likelihood with the sum of squares loss. This implies Theorem 4.1 can be extended to sub-Gaussian edge distributions, and in particular to the RDPG binary edge model with $\Theta_{k\ell}\in[0,1]^{n\times n}$ and $A_{k\ell}\sim\operatorname{Bernoulli}(\Theta_{k\ell}),(k,\ell)\in\mathcal{I}$.
Theorem 4.1 expresses the rates in terms of key quantities such as $r_{k}^{(I)}$, $r^{(II)}$, $m_{k}$, and $n$, allowing us to apply the analysis to a wide range of scenarios. For instance, we can derive the conditions under which Algorithm 1 produces estimates matching the oracle error rates, where each parameter $S,\{Q_{k}\}_{k=1}^{K},\{R_{k\ell}\}_{\mathcal{I}}$ is estimated using its true rank and true values of all other parameters:
$\hat{S}^{(oracle)}=\Bigl[\frac{1}{M}\sum_{\mathcal{I}}\left(A_{k\ell}-Q_{k}-R_{k\ell}\right)\Bigr]_{d_{0}}=\bigl[S+\bar{E}\bigr]_{d_{0}}\ ,$ (23)
$\hat{Q}_{k}^{(oracle)}=\Bigl[\frac{1}{m_{k}}\sum_{\ell=1}^{m_{k}}\left(A_{k\ell}-S-R_{k\ell}\right)\Bigr]_{d_{k}}=\bigl[Q_{k}+\bar{E}_{k}\bigr]_{d_{k}}\ ,$
$\hat{R}_{k\ell}^{(oracle)}=\left[A_{k\ell}-S-Q_{k}\right]_{d_{k\ell}}=\bigl[R_{k\ell}+E_{k\ell}\bigr]_{d_{k\ell}}\ .$
where the second set of equalities follows from (19). Per Lemma B.4, this shows that their errors are dominated by the rates for the corresponding noise components in (20). We state sufficient conditions ensuring that Algorithm 1 achieves the oracle rates in the following corollary of Theorem 4.1. We begin by stating an additional assumption on the rates of group sizes and pairwise similarities of individual and group components.
Assumption 4.
It holds $s_{u,u}\lesssim n^{1/2-\tau},\ s_{w,w}\lesssim K^{-1},$ and $m_{k}\lesssim n^{\tau-1/2}$ for each $k=1,\ldots,K$.
Corollary 4.1 (Oracle rate conditions).
Under Assumption 4 and assumptions of Theorem 4.1 with $r^{(II)}\lesssim(nK/M)^{1/2}$ and $r_{k}^{(I)}\lesssim n^{1/2}$ for each $k=1,\ldots,K$, it holds
$\begin{array}[]{lll}\|\hat{R}_{k\ell}-R_{k\ell}\|_{F}\leq C_{R}d_{k\ell}^{1/2}n^{1/2},&\|\widehat{S+Q}_{k}-(S+Q_{k})\|_{F}\leq C_{S+Q}(d_{0}+d_{k})^{1/2}n^{1/2}m_{k}^{-1/2}\\ \|\hat{Q}_{k}-Q_{k}\|_{F}\leq C_{Q}d_{k}^{1/2}n^{1/2}m_{k}^{-1/2},&\|\hat{S}-S\|_{F}\leq C_{S}d_{0}^{1/2}n^{1/2}M^{-1/2},\end{array}$
if the hyperparameters are set with sufficiently large positive constants $c_{1k}$ and $c_{2}$ as
$\lambda_{1k}=c_{1k}\sqrt{nm_{k}},\quad\alpha_{1k\ell}=1/\sqrt{m_{k}},\quad\lambda_{2}=c_{2}\sqrt{nM},\quad\alpha_{2k}=\sqrt{m_{k}/M},\quad\quad(k,\ell)\in\mathcal{I}.$ (24)
We conclude this section by stating sufficient conditions, which guarantee that our initializers of $\{S+Q_{k}\}_{k=1}^{K}$ in the first stage and of $S$ in the second stage have the rates as in Corollary (4.1).